Please Help!
The product of:
2 1
-3 5
and
3 1 -1
2 4 -2
all in brackets, is what?
Thank you for your time!
: Please Help!
The product of:
2 1
-3 5
and
3 1 -1
2 4 -2
all in brackets, is what?
Thank you for your time!

Answer by Fombitz(1744)

In brackets, I'm assuming mean matrix notation.
A 2x2 matrix multiplied by a 2x3 matrix will produce a 2x3 matrix.
.
.
.
(matrix(2,2,2,1,-3,5))(matrix(2,3,3,1,-1,2,4,-2))=(matrix(2,3,a,b,c,d,e,f))
.
.
.
The rows of the first matrix are multiplied by the columns of the second matrix in this manner,
a=2*3+1*2=8
b=2*1+1*4=6
c=2*(-1)+1*(-2)=-4
d=(-3)*3+5*2=1
e=(-3)*1+5*4=17
f=(-3)*(-1)+5*(-2)=-7
.
.
.
(matrix(2,2,2,1,-3,5))(matrix(2,3,3,1,-1,2,4,-2))=(matrix(2,3,8,6,-4,1,17,-7))